# RTMap: Real-Time Recursive Mapping with Change Detection and Localization

### v1

@@ 1 Introduction
Experiments on localization, mapping, and map change detection tasks on multiple publicly available datasets [17, 3] have demonstrated our effectiveness in simultaneously addressing them through an end-to-end framework. We will release our labeled assets on these datasets and our multi-task model to facilitate future research.
@@ 2.1 Online HD Mapping
Single-traversal temporal fusion methods. To facilitate the stability of online HD mapping between consecutive frames, many subsequent methods [39, 20] leverage BEVFormer [22] for temporal self-attention to fuse temporally adjacent BEV information recurrently. Other explicit methods [6] use odometry to transform perceived results and perform voxel mapping for multi-frame fusion.
Prior-aided methods. Previous works [31, 1] demonstrate that prior-informed online mapping can further improve completeness and stability. Unlike the single-traversal methods mentioned above, prior-aided methods require both localization and change detection to establish correct correspondences between maps. We mainly found two categories of prior-aided methods for online HD mapping: The first category [37, 35] uses neural representation as prior-maps, which requires dense storage unsuitable for mass production, and it is also not intuitive for quality inspection and manual correction. The second category uses Standard-Definition (SD) or HD maps as prior knowledge: SMERF [27] queries the SD map at the ego vehicle’s location for lane-topology reasoning, and U-BEV [4] extracts BEV features from both sensors and SD map to reach meters-level accuracy through template matching. P-MapNet [16] leverages SD priors for coarse localization and HD priors for final map refinement. Our method belongs to the latter category, stepping further to reach real-time centimeter-level localization accuracy beyond previous methods through an end-to-end model. Meanwhile, RTMap manages to self-enclose the maintenance of prior HD maps, inferencing the probabilistic density of online HD map elements for noise-aware HD crowdsourcing.
@@ 3.4 Loss Function for Map Elements and Poses
Composite Geometric Loss for Element Vertices. Following Gu et al. [12], we choose two uni-variate Gaussian distributions to model map elements, allowing us to additionally output the probabilistic density $\mathbf{U}_{t}$ jointly with the location $\mathbf{M}_{t}$ of map element vertices. Specifically, we augment the vertex of each element with two standard deviations on the $\mathrm{XOY}$ plane, regarding them as horizontal 2D Gaussians for obtaining a probabilistic representation. For inferencing such additional information, we augment a negative log-likelihood (NLL) loss $\mathbf{L}_{\mathrm{nll}}$ for each vertex, as:
@@
where $\mathbf{m}|_{\mathrm{xy}}\sim\mathcal{N}^{2}(\mathbf{\mu},\mathbf{\sigma})$ incorporates a 2D Gaussian distribution for each point’s $\mathrm{XOY}$ coordinate. We keep their $\mathrm{Z}$-value representation unchanged due to insufficient observability. The parameters $\mu_{v}^{k},\sigma_{v}^{k}\in\mathbb{R}$ denote the location and scale parameters of the $k^{\mathrm{th}}$ dimensional Gaussian distribution for a vertex $\mathbf{m}_{v}$, and we use a fixed total size $\mathrm{V}=20$ for each map element. We remain the original manhattan distance loss $\mathbf{L}_{\mathrm{mht}}$ [24] with hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ to balance the convergence ability and the element accuracy.
@@ 3.5 MAP for Localization and Crowdsourcing
where $\mathbf{E}^{1}(\mathbf{D}_{t})=\sum_{\mathbf{D}_{t}}-\log(\mathbf{f}^{1})$ is the sum of the negative log-likelihood of point-to-point residuals $\mathbf{f}^{1}$, and we use the squared Mahalanobis distance whose covariance is assigned as the Gaussian mixture $\mathbf{g}_{t}^{i}\triangleq\mathbf{u}_{t}^{i}\oplus u_{t-1}^{i}$ of both inferenced and crowdsourced probabilistic densities [5]. Each pair of associations denoted as $d_{t}^{i}:\left\langle\mathbf{m}_{t}^{i},m_{t-1}^{i}\right\rangle\in\mathbf{D}_{t}$ constructs a residual between corresponded vertices across two maps. One can combine more residuals like IMU pre-integration [11] or odometry [29, 41] for a tightly coupled system, and in this paper, we mainly test the difference between the single residual form $\mathbf{E}^{1}$ and the end-to-end regressed pose $\mathbf{T}^{\mathbb{E}}$.
@@ 4 Implementation, Results, and Evaluation
4.1 Implementation details.
Training Sample Generation. To better align with real-world scenarios and improve the robustness of the network, we first apply random perturbations to the ground truth pose $\mathbf{T}^{\mathrm{gt}}$ to generate a noisy initial pose estimate $\mathbf{T}^{\mathrm{init}}$. Based on $\mathbf{T}^{\mathrm{init}}$, we then crop out the ground truth prior-map. Occasionally, when the local map obtained from the noisy pose is the same size as the perceptual range, some map elements may fall outside the perceptible area. To address this, we add padding to the perceptual range to ensure that all elements of the ground truth map remain within the perceptible region. Following this, we randomly remove certain elements from the prior-map and generate synthetic fake map elements. We do not perturb the ground truth map elements to simulate specific changes (such as shifting lane markings), as such modifications could be interpreted as deleting the original map elements and adding new ones.
Implementation and Training Details. We choose ResNet50 [13] as the image backbone. Our RTMap model is trained using 8 NVIDIA GeForce RTX 3090 GPUs with a total batch size of 32 for 36 epochs. We utilize the AdamW optimizer with a learning rate of $6\times 10^{-4}$. To ensure fair comparisons, we adopt the same training settings for MapTRv2 [24] and ensure its convergence. For $\mathbf{L}_{\mathrm{pts}}$ in Eq. 3, we set $\lambda_{1}=0.03$ and $\lambda_{2}=5.0$ for balancing their influence. For $\mathbf{L}_{\mathrm{pose}}$, the translation weight is set to $0.04$, corresponding to the scale in radians used when predicting the heading angle. We set the longitudinal range to $[-36,36]$ and the lateral range to $[-18,18]$ meters for online HD mapping.
@@ 4.2 Experiment Setup
Tasks and Metrics. Regarding the sufficiency of multi-traversal scanning and map-changing events, we test all sub-tasks, including crowdsourcing, change detection, and localization on the TbV dataset. Meanwhile, we use the nuScenes dataset to test the performance of our map-based localization w.r.t. its labeled HD map.
For the crowdsourced mapping task, we maintain the evaluation method used in online HD mapping tasks to assess the mapping results compared to single-traversal methods. To obtain ground truth for evaluation, we augment the map data by aligning the previous validation set with sensor data for ensuring 3D consistency. Specifically, we use average precision to evaluate the quality of the generated map and Chamfer distance to determine the alignment between the ground truth and predicted maps. We compute the AP at the following thresholds $\{0.5,1.0,1.5\}$ in meters keeping the same with experiments in other methods [24], and then calculate the mean to obtain the final mean average precision ($\mathrm{mAP}$).
For the change detection task, we follow the evaluation implementation of the officially released codehttps://github.com/johnwlambert/tbv in TbV [17], to treat the change detection task as a binary classification problem ($\mathrm{Acc}$), with classes indicating whether a change has occurred $\mathrm{Acc}_{c}$ or not $\mathrm{Acc}_{r}$, and the mean accuracy ($\mathrm{mAcc}$) is also calculated to assess overall performance.
@@ 4.3 Performance on Crowdsourcing
We cluster multi-traversal spatially overlapped driving clips in TbV for performing crowdsourcing, group 15 found clips into two scenarios – straight (6) and turning (9), and listed our results in comparison to our baseline MapTRv2 [24] as in Tab. 1. While single-traversal methods perform consistently among different cycles, our crowdsourced mapping framework gradually improves the prior-aided map quality through multiple cycles. We also compare to a direct version, which does not infer and use probabilistic densities $\mathbf{U}$ during crowdsourcing, highlighting the effectiveness of leveraging such uncertainty in mapping tasks. We also refer readers to Fig. 4, which presents the trend of quality improvement by solving a precise trajectory and experiencing multi-traversals.
@@ 5 Conclusion
Future work. Our future directions include incorporating more sensors (e.g., LiDAR) to form a multi-modal sensor fusion and uploading auxiliary but highly compact data for better probabilistic modeling on the cloud. Expanding validation to unstructured terrains and more complex urban environments will better establish the framework’s generalizability.
@@ Acknowledgements
We thank the reviewers for the valuable discussions. We respect to our past and present colleagues for their unwavering dedication to pursuing innovative lighter-map solutions, whose contributions have profoundly bring-up this work. This research was supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F030001.

### v2

@@ 1 Introduction
Experiments on localization, mapping, and map change detection tasks on multiple publicly available datasets [18, 3] have demonstrated our effectiveness in simultaneously addressing them through an end-to-end framework.
@@ 2.1 Online HD Mapping
Single-traversal temporal fusion methods. To facilitate the stability of online HD mapping between consecutive frames, many subsequent methods [41, 21] leverage BEVFormer [23] for temporal self-attention to fuse temporally adjacent BEV information recurrently. Other explicit methods [7] use odometry to transform perceived results and perform voxel mapping for multi-frame fusion. MapTracker [6] leverages BEV and vector memory fusion to better exploit temporal information. PrevPredMap [30] use previous predictions as priors.
Prior-aided methods. Previous works [33, 1] demonstrate that prior-informed online mapping can further improve completeness and stability. Unlike the single-traversal methods mentioned above, prior-aided methods require both localization and change detection to establish correct correspondences between maps. We mainly found two categories of prior-aided methods for online HD mapping: The first category [39, 37] uses neural representation as prior-maps, which is storage-intensive and not suitable for large-scale deployment or manual inspection. The second category uses Standard-Definition (SD) or HD maps as prior knowledge: SMERF [28] queries the SD map at the ego vehicle’s location for lane-topology reasoning, and U-BEV [4] extracts BEV features from both sensors and SD map to reach meters-level accuracy through template matching. P-MapNet [17] leverages SD priors for coarse localization and HD priors for final map refinement. PriorDrive [42] integrates diverse prior maps through the Unified Vector Encoder. Our method belongs to the latter category, stepping further to reach real-time centimeter-level localization accuracy beyond previous methods through an end-to-end model. Meanwhile, RTMap manages to self-enclose the maintenance of prior HD maps, inferencing the probabilistic density of online HD map elements for noise-aware HD crowdsourcing.
@@ 3.4 Loss Function for Map Elements and Poses
Composite Geometric Loss for Element Vertices. Following Gu et al. [13], we choose univariate Laplace distributions to model map elements, allowing us to additionally output the uncertainty of the location $\mathbf{M}_{t}$ of map element vertices. For inferencing such additional information, we augment a negative log-likelihood (NLL) loss $\mathbf{L}_{\mathrm{nll}}$ for each vertex, as:
@@
We keep their $\mathrm{Z}$-value representation unchanged due to insufficient observability. The parameters $\mu_{v}^{k},\sigma_{v}^{k}\in\mathbb{R}$ denote the location and scale parameters of the $k^{\mathrm{th}}$ dimensional Laplace distributions for a vertex $\mathbf{m}_{v}$, and we use a fixed total size $\mathrm{V}=20$ for each map element. We remain the original manhattan distance loss $\mathbf{L}_{\mathrm{mht}}$ [25] with hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ to balance the convergence ability and the element accuracy.
@@ 3.5 MAP for Localization and Crowdsourcing
where $\mathbf{E}^{1}(\mathbf{D}_{t})=\sum_{\mathbf{D}_{t}}-\log(\mathbf{f}^{1})$ is the sum of the negative log-likelihood of point-to-point residuals $\mathbf{f}^{1}$, and we use the squared Mahalanobis distance whose covariance is assigned as the Gaussian mixture $\mathbf{g}_{t}^{i}\triangleq\mathbf{u}_{t}^{i}\oplus u_{t-1}^{i}$ of both inferenced and crowdsourced probabilistic densities [5]. Each pair of associations denoted as $d_{t}^{i}:\left\langle\mathbf{m}_{t}^{i},m_{t-1}^{i}\right\rangle\in\mathbf{D}_{t}$ constructs a residual between corresponded vertices across two maps. One can combine more residuals like IMU pre-integration [12] or odometry [31, 45] for a tightly coupled system, and in this paper, we mainly test the difference between the single residual form $\mathbf{E}^{1}$ and the end-to-end regressed pose $\mathbf{T}^{\mathbb{E}}$. Specifically, we use the Levenberg-Marquardt algorithm to minimize $\mathbf{E}^{1}(\mathbf{D}_{t})$.
@@ 4 Implementation, Results, and Evaluation
4.1 Implementation details
Training Sample Generation. To better align with real-world scenarios and enhance network robustness, we first apply random perturbations to the ground truth pose $\mathbf{T}^{\mathrm{gt}}$ to generate a noisy initial pose estimate $\mathbf{T}^{\mathrm{init}}$. Using $\mathbf{T}^{\mathrm{init}}$, we then crop the ground truth map. Occasionally, when the local map obtained from the noisy pose is the same size as the perceptual range, some map elements may fall outside the perceptible area. To ensure all ground truth map elements remain accessible, we pad the perceptual range as necessary. Next, we randomly remove certain elements from the prior-map and generate synthetic fake map elements. We do not perturb the ground truth map elements to simulate specific changes (such as shifting lane markings), as such modifications could be interpreted as deleting the original map elements and adding new ones.
Implementation and Training Details. We choose ResNet50 [14] as the image backbone. Our RTMap model is trained using 8 NVIDIA GeForce RTX 3090 GPUs with a total batch size of 32 for 36 epochs. We utilize the AdamW optimizer with a learning rate of $6\times 10^{-4}$. To ensure fair comparisons, we adopt the same training settings for MapTRv2 [25] and ensure its convergence. For $\mathbf{L}_{\mathrm{pts}}$ in Eq. 3, we set $\lambda_{1}=0.03$ and $\lambda_{2}=5.0$ for balancing their influence. For $\mathbf{L}_{\mathrm{pose}}$, the translation weight is set to 0.040.040.04, corresponding to the scale in radians used when predicting the heading angle. We set the longitudinal range to [36,36][-36,36][ - 36 , 36 ] and the lateral range to [18,18][-18,18][ - 18 , 18 ] meters for online HD mapping. In contrast to prior-aided method such as HRMapNet [43], we do not use crowdsourced prior-maps for training. Instead, we use artificially perturbed prior-maps to ensure we have incorporated sufficient map-changing events regarding the current scale and maturity of publicly available datasets.
@@ 4.2 Experiment Setup
Tasks and Metrics. Regarding the insufficiency of multi-traversal scanning and map-changing events, we tested the main task – crowdsourcing – as well as two sub-tasks, change detection and localization, on the TbV dataset. Meanwhile, we use the nuScenes dataset to test the performance of our map-based localization w.r.t. its labeled HD map.
[…]
For the change detection task, we follow the evaluation implementation of the officially released code in TbV [18], to treat the change detection task as a binary classification problem ($\mathrm{Acc}$), with classes indicating whether a change has occurred $\mathrm{Acc}_{c}$ or not $\mathrm{Acc}_{r}$, and the mean accuracy ($\mathrm{mAcc}$) is also calculated to assess overall performance.
@@ 4.3 Performance on Crowdsourcing
We cluster multi-traversal spatially overlapped driving clips in TbV for performing crowdsourcing, group 15 found clips into two scenarios – straight (6) and turning (9), and listed our results in comparison with state-of-the-art online HD mapping methods, as shown in Tab. 4.3. Due to the introduction of localization noise, the accuracy of existing online HD mapping methods is significantly reduced. Our approach is essentially the same as existing methods in the first cycle, when neither localization nor change detection is required. In later cycles, by utilizing the prior-map, our method gradually decreases localization errors and improves mapping accuracy. We also compare to a direct version, which does not infer and use probabilistic densities $\mathbf{U}$ during crowdsourcing, highlighting the effectiveness of leveraging such uncertainty in mapping tasks. We also refer readers to Fig. 5, which presents the trend of quality improvement by solving a precise trajectory and experiencing multi-traversals.
@@ 5 Conclusion
Future work. Our future directions include incorporating more sensors (e.g., LiDAR) to form a multi-modal sensor fusion and uploading auxiliary but highly compact data for better probabilistic modeling on the cloud. Expanding validation to unstructured terrains and more complex urban environments will better establish the framework’s generalizability. Besides, crowdsourcing enables a maximum-a-posteriori policy to vote out temporary results gradually in our current version. In the future, our backbone supports adding occlusion detection head as an auxiliary task to distinguish between permanent structural changes and transient obstructions.
@@ Acknowledgements
We thank the reviewers for the valuable discussions. We respect our past and present colleagues for their unwavering dedication to pursuing innovative light-weight map solutions, whose contributions have profoundly initiated this work. This research was supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F030001.